• Log InLog In
  • Register
Liquid`
Team Liquid Liquipedia
EDT 16:04
CEST 22:04
KST 05:04
  • Home
  • Forum
  • Calendar
  • Streams
  • Liquipedia
  • Features
  • Store
  • EPT
  • TL+
  • StarCraft 2
  • Brood War
  • Smash
  • Heroes
  • Counter-Strike
  • Overwatch
  • Liquibet
  • Fantasy StarCraft
  • TLPD
  • StarCraft 2
  • Brood War
  • Blogs
Forum Sidebar
Events/Features
News
Featured News
[ASL22] Finals Preview: Sweet Victory!9[ASL22] Ro4 Preview: Mirror Mirror6[ASL22] Ro8 Preview: Within Reach5[ASL22] Ro8 Preview: In A Tizzy11[ASL22] Ro16 Preview: Holy Diver5
Community News
ISSL (IntoTheTV x SOOP SC2 League): Premier2Weekly Cups (October 12-18): 5.0.17 era begins1[$1250] International ShieldBattery League49Weekly Cups (October 5-11): Farewell to 5.0.164Rollback Netcode in Brood War144
StarCraft 2
General
How do you feel about the mass reverts in the 5.0.17 PTR? No more SC2 streamers on team liquid? Weekly Cups (October 12-18): 5.0.17 era begins Server Blocker SCII needs to put more weight on strategy and tact
Tourneys
WORTEX Challenger Series — Open SC2 LAN in Budapes RSL goes to London! 2026 Offline Finals Nov 21-22 Sparkling Tuna Cup - Weekly Open Tournament Stellar Fest TWO the Moon (Dec 16-20) Master Swan Open (Global Bronze-Master 2)
Strategy
[H] ZvP Mid-Late Game: Stalkers Collossi HT
Custom Maps
[G] How to playtest arcade maps [A] Nemrods 1/4 players [M] (2) Sweltering Sands [M] (2) Frigid Storage
External Content
Mutation # 548 Shining Bright The PondCast: SC2 News & Results Mutation # 547 Temple Of Terror Mutation # 546 Catch the Train
Brood War
General
Do we even care about balance? SCHNAIL now supports Remastered BW General Discussion SOOP Announces New Women's Monthly Tournament Mineral gathering rate (boosted) for each patch
Tourneys
[ASL22] Grand Finals [$1250] International ShieldBattery League [Women's StarCraft] Tier 8 CK 2026-10-20 [Megathread] Daily Proleagues
Strategy
Cliff Jump Revisited (1 in a 1000 strategy) Simple Questions, Simple Answers Replay Review Process - What do you do?
Other Games
General Games
Stormgate/Frost Giant Megathread Nintendo Switch Thread Armies of Exigo - YesYes? Dawn of War IV Warcraft III: The Frozen Throne
Dota 2
Dota 2 Champions League Season 3 Begins April 25!
League of Legends
[TL LoL EUW IHs] Teemo shall perish
Heroes of the Storm
Heroes of the Storm 2.0
Hearthstone
Deck construction bug
TL Mafia
TL Mafia Community Thread
Community
General
Canadian Politics Mega-thread US Politics Mega-thread The Chess Thread Things Aren’t Peaceful in Palestine Artificial Intelligence Thread
Fan Clubs
Serral Fan Club
Media & Entertainment
Movie Discussion! [Manga] One Piece
Sports
TeamLiquid Health and Fitness Initiative For 2023 MLB/Baseball 2023 Football (Soccer) Thread
World Cup 2022
Tech Support
Computer Build, Upgrade & Buying Resource Thread
TL Community
Recent Gifted Posts
Blogs
How Much Genes Contribute to…
TrAiDoS
ShieldBattery on Mac
Kraekkling
[ASL22] Ro4 Day1 Ticket Giv…
bITt.mAN
38 yo Retired SWE loo…
PurE)Rabbit-SF
Can Bots Beat Pros?? Starcr…
namkraft
Customize Sidebar...

Website Feedback

Closed Threads



Active: 10460 users

Probability

Blogs > paper
Post a Reply
paper
Profile Blog Joined September 2004
13197 Posts
Last Edited: 2010-12-03 07:42:23
February 19 2009 05:50 GMT
#1
[image loading]


+ Show Spoiler +
yoyoyo! i have a midterm tomorrow and i have some problems that i don't have the solutions to!

if you're bored, you can check my work, or if you just love probabilities that much, you can try to do them yourself without checking the solns! :}

[all my XC0 will be for clarity :d]



P9. 10 rooms in a ring. Each of the 3 Little Pigs checks into a different room. Let X = # of adjacent pairs. Find P(X>=1).
[spoiler=My Solution]P(X>=1) = 1-P(X=0)
P(X=0) = 15/10C3
P(X>=1) = 105/120[/spoiler]

P10. Draw 10 cards from standard deck. Find:
a) P(getting exactly 1 suit)
b) P(getting exactly 2 suits)
[spoiler=My Solution]
10a. 4C1*13C10*39C0 / 52C10
10b. 4C2*[13C1*13C9+13C2*13C8+...+13C9*13C1]*26C0 / 52C10
[/spoiler]

P14. Return 4 hats randomly back to 4 people (n=4). Let X=# matches. Find:
a. P(X=4)
b. P(X=3)
c. P(X>=1)
d. P(X=-)
[spoiler=My Solution]
a. 1/4P4 = 1/24
b. if you have 3 matching, you must also have 4 matching, so same as (a)? is there overlap?
c. 1-(6/24)
d. 6/24[/spoiler]

will add more as i continue doing practice problems...^^


***
Hates Fun🤔
Abydos1
Profile Blog Joined October 2008
United States832 Posts
February 19 2009 06:12 GMT
#2
On February 19 2009 14:50 paper wrote:
P14. Return 4 hats randomly back to 4 people (n=4). Let X=# matches. Find:
a. P(X=4)
b. P(X=3)
c. P(X>=1)
d. P(X=-)
+ Show Spoiler [My Solution] +

a. 1/4P4 = 1/24
b. if you have 3 matching, you must also have 4 matching, so same as (a)? is there overlap?
c. 1-(6/24)
d. 6/24


will add more as i continue doing practice problems...^^


+ Show Spoiler [part b] +

Actually, you're on the right track for this one but wrong. As you said if you have 3 matching the 4th must match too therefore P(X=3) = 0. You can never have exactly 3 pairs matching.
"...perhaps the greatest joy possible in Starcraft, being accused of being a maphacker" - Day[9]
OmgIRok
Profile Blog Joined June 2008
Taiwan2699 Posts
February 19 2009 06:13 GMT
#3
Wait, what was that picture for?

Funny how I had a blog today with help on probability/statistics.
I don't think you posted in there to help me though haha x_X
"Wanna join my [combo] clan?" "We play turret d competitively"
Cambium
Profile Blog Joined June 2004
United States16368 Posts
Last Edited: 2009-02-19 06:16:41
February 19 2009 06:16 GMT
#4
nice pic ^^

I thought it was going to be along the lines of "what's the probability of me banging her at the back of the club next time i see her"
When you want something, all the universe conspires in helping you to achieve it.
paper
Profile Blog Joined September 2004
13197 Posts
Last Edited: 2009-02-19 06:17:34
February 19 2009 06:17 GMT
#5
On February 19 2009 15:13 OmgIRok wrote:
Wait, what was that picture for?

Funny how I had a blog today with help on probability/statistics.
I don't think you posted in there to help me though haha x_X



hahah i actually did your problems on my scratch paper but 93819023 people answered before i had even started on them anyway :s

btw thanks abydos
Hates Fun🤔
Abydos1
Profile Blog Joined October 2008
United States832 Posts
February 19 2009 06:20 GMT
#6
On February 19 2009 15:16 Cambium wrote:
nice pic ^^

I thought it was going to be along the lines of "what's the probability of me banging her at the back of the club next time i see her"


Lol, the pic didn't even load for me the first time I viewed this.
"...perhaps the greatest joy possible in Starcraft, being accused of being a maphacker" - Day[9]
Cambium
Profile Blog Joined June 2004
United States16368 Posts
February 19 2009 06:38 GMT
#7
Here's a better probability question:

You have a regular fair die and you are allowed three rolls, and you can stop whenever you want. I will give you X dollars at the end where X is the face value of the die. Find the EV.
When you want something, all the universe conspires in helping you to achieve it.
SayaSP
Profile Blog Joined February 2007
Laos5494 Posts
February 19 2009 07:21 GMT
#8
Hey paper! I can't answer ur Questionz but hi!
[iHs]SSP | I-NO-KI BOM-BA-YE | のヮの http://tinyurl.com/MLIStheCV , MLIS.
SpiritoftheTunA
Profile Blog Joined August 2006
United States20903 Posts
February 19 2009 07:24 GMT
#9
On February 19 2009 15:38 Cambium wrote:
Here's a better probability question:

You have a regular fair die and you are allowed three rolls, and you can stop whenever you want. I will give you X dollars at the end where X is the face value of the die. Find the EV.

doesn't the value a person would generally stop at change the EV for them? oh well

its probably around seventy
posting on liquid sites in current year
OmgIRok
Profile Blog Joined June 2008
Taiwan2699 Posts
February 19 2009 07:27 GMT
#10
On February 19 2009 15:17 paper wrote:
Show nested quote +
On February 19 2009 15:13 OmgIRok wrote:
Wait, what was that picture for?

Funny how I had a blog today with help on probability/statistics.
I don't think you posted in there to help me though haha x_X



hahah i actually did your problems on my scratch paper but 93819023 people answered before i had even started on them anyway :s

btw thanks abydos


ohoh haha nice i'll refer to you next time i need help with this probability/stats you look like the expert :O

So what was that picture for?
"Wanna join my [combo] clan?" "We play turret d competitively"
SpiritoftheTunA
Profile Blog Joined August 2006
United States20903 Posts
February 19 2009 07:29 GMT
#11
On February 19 2009 16:27 OmgIRok wrote:
Show nested quote +
On February 19 2009 15:17 paper wrote:
On February 19 2009 15:13 OmgIRok wrote:
Wait, what was that picture for?

Funny how I had a blog today with help on probability/statistics.
I don't think you posted in there to help me though haha x_X



hahah i actually did your problems on my scratch paper but 93819023 people answered before i had even started on them anyway :s

btw thanks abydos


ohoh haha nice i'll refer to you next time i need help with this probability/stats you look like the expert :O

So what was that picture for?

to unexpectedly brighten your day!
posting on liquid sites in current year
TheYango
Profile Joined September 2008
United States47024 Posts
Last Edited: 2009-02-19 07:40:28
February 19 2009 07:40 GMT
#12
On February 19 2009 16:24 SpiritoftheTunA wrote:
Show nested quote +
On February 19 2009 15:38 Cambium wrote:
Here's a better probability question:

You have a regular fair die and you are allowed three rolls, and you can stop whenever you want. I will give you X dollars at the end where X is the face value of the die. Find the EV.

doesn't the value a person would generally stop at change the EV for them? oh well

its probably around seventy

I assume the problem implies that a person will always make the statistically favorable choice, e.g. stop when the expected value of the remaining rolls is lower than or equal to their current roll.

+ Show Spoiler +

Working backwards:
EV of 1 roll = 7/2
For 2 rolls, keep roll on 4, 5, or 6, re-roll on 1, 2 or 3. EV is therefore:
1/2 * 5 + 1/2 * 7/2 = 17/4
For 3 rolls, keep roll on 5 or 6, re-roll on 1, 2, 3, or 4. EV is therefore:
2/3 * 4.25 + 1/3 * 5.5 = 14/3
Moderator
Cambium
Profile Blog Joined June 2004
United States16368 Posts
Last Edited: 2009-02-19 08:24:04
February 19 2009 08:16 GMT
#13
On February 19 2009 16:40 TheYango wrote:
Show nested quote +
On February 19 2009 16:24 SpiritoftheTunA wrote:
On February 19 2009 15:38 Cambium wrote:
Here's a better probability question:

You have a regular fair die and you are allowed three rolls, and you can stop whenever you want. I will give you X dollars at the end where X is the face value of the die. Find the EV.

doesn't the value a person would generally stop at change the EV for them? oh well

its probably around seventy

I assume the problem implies that a person will always make the statistically favorable choice, e.g. stop when the expected value of the remaining rolls is lower than or equal to their current roll.

+ Show Spoiler +

Working backwards:
EV of 1 roll = 7/2
For 2 rolls, keep roll on 4, 5, or 6, re-roll on 1, 2 or 3. EV is therefore:
1/2 * 5 + 1/2 * 7/2 = 17/4
For 3 rolls, keep roll on 5 or 6, re-roll on 1, 2, 3, or 4. EV is therefore:
2/3 * 4.25 + 1/3 * 5.5 = 14/3


For those that are interested, this was given to me during a Google interview. I gave the "correct" answer, which, later, after discussing with friends, I realized was actually probably incorrect (I haven't gotten around to work it out yet).

I said exactly what you said...
+ Show Spoiler +

I think that's incorrect because you should stop only when you roll is lower than your future potential EV.

EV of 1 roll = 7/2
So, technically, you should re-roll if you get 1, 2 or 3.

However, if you look at it from the other end, you still have two rolls left, which will yield an EV of 17/4, so I think you should actually stop rolling only if you get 5 or 6 on your first roll.


Okay, I just tried to work it out with what I said in the spoiler, and I just cannot put it to work as I got really confused with forward and backward EV. (but in my defense, it's 3:30 AM), let me know if you figure it out, or if it doesn't make sense at all.
When you want something, all the universe conspires in helping you to achieve it.
TheYango
Profile Joined September 2008
United States47024 Posts
Last Edited: 2009-02-19 17:48:32
February 19 2009 17:44 GMT
#14
On February 19 2009 17:16 Cambium wrote:
+ Show Spoiler +

I think that's incorrect because you should stop only when you roll is lower than your future potential EV.

EV of 1 roll = 7/2
So, technically, you should re-roll if you get 1, 2 or 3.

However, if you look at it from the other end, you still have two rolls left, which will yield an EV of 17/4, so I think you should actually stop rolling only if you get 5 or 6 on your first roll.

+ Show Spoiler +

Um, you did the exact same thing I did.

Your first roll keeps 5s or 6s, meaning that the EV of all 1st rolls that are kept is 5.5. If you choose to re-roll, your second roll is kept on 4, 5, or 6, meaning the EV of all 2nd rolls that are kept is 5. If you roll a 3rd time, you have an EV of 3.5.
So the total EV is:

(1/3)*5.5 + (2/3)*(1/2)*5 + (2/3)*(1/2)*3.5 = 14/3
Moderator
Please log in or register to reply.
Live Events Refresh
Next event in 4h 26m
[ Submit Event ]
Live Streams
Refresh
StarCraft 2
mouzHeroMarine 538
UpATreeSC 103
BRAT_OK 65
JuggernautJason52
StarCraft: Brood War
Shuttle 649
ProTech162
Mind 90
Dota 2
syndereN345
Counter-Strike
fl0m1540
Fnx 1096
Super Smash Bros
C9.Mang0172
Heroes of the Storm
Grubby3978
Liquid`Hasu463
Trikslyr97
Other Games
summit1g2818
FrodaN956
Beastyqt772
B2W.Neo503
Pyrionflax238
ZombieGrub5
Railgan3
Organizations
Other Games
WardiTV244
[ Show 16 non-featured ]
StarCraft 2
• Hinosc 34
• Reevou 5
• AfreecaTV YouTube
• intothetv
• Kozan
• IndyKCrew
• Migwel
StarCraft: Brood War
• 80smullet 33
• blackmanpl 22
• BSLYoutube
• STPLYoutube
• ZZZeroYoutube
League of Legends
• TFBlade458
Other Games
• imaqtpie1155
• Shiphtur626
• WagamamaTV378
Upcoming Events
PiGosaur Cup
4h 26m
The PondCast
13h 56m
Kung Fu Cup
14h 56m
IntoTheTV X SOOP
1d 14h
Reynor vs Zoun
Neeb vs Rogue
FEL
3 days
Patches Events
3 days
AI Arena Tournament
3 days
FEL
4 days
Sparkling Tuna Cup
4 days
Replay Cast
5 days
[ Show More ]
WardiTV Weekly
5 days
IntoTheTV X SOOP
6 days
ByuN vs Lambo
MaxPax vs Bunny
Liquipedia Results

Completed

ASL Season 22
Blizzard Classic Cup 2026
HCC Season 3

Ongoing

Acropolis #5
IPSL Winter 2026-27
KCM Race Survival 2026 Season 4
Acropolis #5 - GSC
1win Private Club #2
Thunderpick World Champ. '26
ESL Pro League Season 24
Stake Ranked Episode 4
1win Private Club #1
Logitech G Play Connect 2026
SL StarSeries Fall 2026
FISSURE Playground #3
BLAST Open Fall 2026
Esports World Cup 2026

Upcoming

BSL Season 23
SC4ALL II: Brood War
BSL 23: Non-Korean Championship
HSC XXX
Stellar Fest 2: Lunar Cup
SC4ALL II: StarCraft II
Kung Fu Cup 2026 Grand Finals
RSL Offline Finals
eXTREMESLAND 2026
PGL Major Singapore 2026
Stake Ranked Episode 6
BLAST Rivals Fall 2026
IEM Beijing 2026
Stake Ranked Episode 5
PGL Masters Bucharest 2026
TLPD

1. ByuN
2. TY
3. Dark
4. Solar
5. Stats
6. Nerchio
7. sOs
8. soO
9. INnoVation
10. Elazer
1. Rain
2. Flash
3. EffOrt
4. Last
5. Bisu
6. Soulkey
7. Mini
8. Sharp
Sidebar Settings...

Advertising | Privacy Policy | Terms Of Use | Contact Us

Original banner artwork: Jim Warren
The contents of this webpage are copyright © 2026 TLnet. All Rights Reserved.